# Thermodynamic Curvature and the Widom Ridge in Interacting Spin Systems

### 17 Apr 2026 v1

@@ II Geometric Formulation of Thermodynamic Response
The work performed over a quasistatic cycle admits a natural geometric interpretation in terms of holonomy. Writing the work as
@@ IV.2 Curvature field and Widom ridge
The resulting curvature field $\phi_{\beta}(\beta,h)\equiv\Omega_{\beta h}$ is shown in Fig. 1. Several key features are immediately apparent.
The field is evaluated using the fluctuation-based estimator $\phi_{\beta}=-N\left(\langle me\rangle-\langle m\rangle\langle e\rangle\right)$, where $m$ and $e$ are the intensive magnetization and energy densities. The color scale indicates the magnitude and sign of the curvature, with a pronounced negative ridge extending from the critical point at $(\beta_{c},0)$ into the supercritical regime. The solid line denotes the Widom ridge, defined as the locus $\beta^{*}(h)=\arg\min_{\beta}\phi_{\beta}(\beta,h)$. The ridge closely follows the region of maximal curvature magnitude, demonstrating that the Widom line corresponds to a geometric feature of the control manifold. As $h\to 0$, the curvature becomes increasingly localized and the ridge sharpens, reflecting the enhancement of correlated energy–magnetization fluctuations near criticality. The gray points indicate the sampling locations used in the Monte Carlo evaluation, with increased density in the vicinity of the ridge to resolve the rapidly varying curvature. This figure provides direct numerical evidence that thermodynamic crossover behavior in the Ising model is governed by the structure of the curvature field.
First, the curvature field exhibits a pronounced negative ridge extending from the critical point at $(\beta_{c},h=0)$ into the supercritical regime. This ridge corresponds to the region where the covariance between energy and magnetization fluctuations is maximized in magnitude, indicating strong coupling between these observables. Second, the curvature becomes increasingly localized as $h\to 0$, reflecting the divergence of thermodynamic response near the critical point. Away from this region, the curvature field varies smoothly and approaches small values, indicating weak coupling between energy and magnetization fluctuations.

### 28 Apr 2026 v2

@@ II Geometric Formulation of Thermodynamic Response
The work performed over a quasistatic cycle admits a natural geometric interpretation in terms of holonomy. i.e., the accumulated change obtained by transporting the system around a closed loop in parameter space, which depends on the path taken rather than only its endpoints. Writing the work as
@@
Relation to geometric response and Berry curvature.
Geometric response functions analogous to Berry curvature have been identified in both quantum and classical nonequilibrium systems, where cyclic driving in parameter space leads to geometric phases and modified response properties [1]. In such settings, the curvature is associated with the breakdown of equilibrium thermodynamic relations and the emergence of geometric contributions to response.
In contrast, the present formulation arises within equilibrium statistical mechanics. Here, the curvature is derived directly from thermodynamic potentials and can be expressed in terms of equilibrium fluctuations. Rather than signaling a departure from thermodynamic structure, it provides a local description of thermodynamic response and encodes the coupling between control variables.
@@ IV.2 Curvature field and Widom ridge
The resulting curvature field $\phi_{\beta}(\beta,h)\equiv\Omega_{\beta h}$ is shown in Fig. 1. The field is evaluated using the fluctuation-based estimator $\phi_{\beta}=-N\left(\langle me\rangle-\langle m\rangle\langle e\rangle\right)$, where $m$ and $e$ are the intensive magnetization and energy densities. The color scale indicates the magnitude and sign of the curvature, with a pronounced negative ridge extending from the critical point at $(\beta_{c},0)$ into the supercritical regime. The solid line denotes the Widom ridge, defined as the locus $\beta^{*}(h)=\arg\min_{\beta}\phi_{\beta}(\beta,h)$. The ridge closely follows the region of maximal curvature magnitude, demonstrating that the Widom line corresponds to a geometric feature of the control manifold. As $h\to 0$, the curvature becomes increasingly localized and the ridge sharpens, reflecting the enhancement of correlated energy–magnetization fluctuations near criticality. The gray points indicate the sampling locations used in the Monte Carlo evaluation, with increased density in the vicinity of the ridge to resolve the rapidly varying curvature. This figure provides direct numerical evidence that thermodynamic crossover behavior in the Ising model is governed by the structure of the curvature field.
Several key features are immediately apparent:
  • First, the curvature field exhibits a pronounced negative ridge extending from the critical point at $(\beta_{c},h=0)$ into the supercritical regime. This ridge corresponds to the region where the covariance between energy and magnetization fluctuations is maximized in magnitude, indicating strong coupling between these observables.
  • Second, the curvature becomes increasingly localized as $h\to 0$, reflecting the divergence of thermodynamic response near the critical point. Away from this region, the curvature field varies smoothly and approaches small values, indicating weak coupling between energy and magnetization fluctuations.